Evergreen Corporation manufactures circuit boards and is in the process of preparing next year's budget. The pro forma income statement for the current year is presented below.

Sales $ 3,900,000
Cost of sales:
Direct material $ 540,000
Direct labor 450,000
Variable overhead 295,000
Fixed overhead 800,000 2,085,000
Gross profit $ 1,815,000
Selling and General & Admin. Exp.
Variable 790,000
Fixed 290,000 1,080,000
Operating income $ 735,000


The contribution margin ratio for the current year is:
Answer #1

Total Variable costs

= 540,000+450,000+295,000+790,000 = 2,075,000

Contribution margin ratio

= (Selling price - Total Variable costs) /Selling price

= (3,900,000-2,075,000)/3,900,000

= 46.8%